Circular No. 6762

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997di IN UGC 4015 Y. L. Qiu, Q. Y. Qiao, W. D. Li, and J. Y. Hu, Beijing Astronom-

ical Observatory (BAO); and A. Esamdin, Urumqi Astronomical Station, on behalf of the BAO supernova survey, report the discovery of a supernova (mag ∼ 18.1) on an unfiltered CCD frame taken with the BAO 0.6-m reflector on Oct. 27.9 UT. SN 1997di is located at α = 7^{h}47^{m}44^{s}.45, δ = +62^{o}19′39′′.2 (equinox 2000.0), which is 29′′ west and 11′′ north of the nucleus of UGC 4015. The supernova was confirmed on images obtained on Oct. 26, at mag ∼ 18.5. Unfiltered CCD images of the same field taken on Oct. 19 (limiting mag ∼ 19.0) show no star at the position of the supernova. A low-dispersion spectrum taken on Oct. 29.8 with the 2.16-m telescope at Xinglong station shows this to be a type-II supernova near maximum. Hα and Hβ display conspicuous P-Cyg profiles. Expansion velocities derived from absorption minimum of Hα and Hβ are 6350 and 5600 km/s, respectively. The recession velocity of the host galaxy is 5620 km/s, as derived from the narrow Hα emission line.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1997di taken by P. Berlind on Oct. 29.5 UT at the Whipple Observatory 1.5-m telescope reveals that it is a normal type-II supernova near maximum, showing P-Cyg Balmer lines. The expansion velocity measured from the Hα profile is 7500 km/s.

SUPERNOVAE The EROS Team (cf. IAUC 6605, 6628, 6760) reports their discovery of

four apparent supernovae on images obtained with the 1-m Marly telescope at the European Southern Observatory:

SN  1997 UT  α_{2000}  δ_{2000}  V  Offset
1997dj Oct. 24 1^{h}14^{m}40.^{s}97 −40^{o}35′51′′.1 22  3′′.6 E, 3′′.0 S
1997dk Oct. 26 23 43 38.69 −41 19 25.4 18.7 8′′.4 W, 7′′.2 S
1997dl Oct. 26 23 43 38.48 −41 19 20.0 19.5 10′′.8 W, 1′′.8 S
1997dm Oct. 26 0 30 12.83 −41 06 08.9 20.3 5′′.7 E, 2′′.7 S

SNe 1997dk and 1997dl are in the same galaxy. SN 1997dj is apparently still rising in brightness. Previous images of these galaxies obtained on 1996 Nov. 1 show no stars at the offsets noted above (limiting mag ∼ 22). All four supernova candidates were confirmed on images taken on Oct. 27.

1997 October 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green